UT 20130901 Technical Observing - Post-Shutdown Wake-up

Observer: JHill, DThompson (Tucson)
Telescope Operator: GBechetti (LBTO)
Software Support: DCox, DSummers, KSummers (Tucson)
AO Support: RBrigulio (LBTO) DMiller (Tucson)
Telescope Support: MMidkiff (LBTO)
Instrument Support: JPowers, ESolheid (LBTO)

Summary

Weather has remarkably improved since Friday, but we are still closed for humidity at the start of the night. We started observing at 4UT after some problems with the shutter doors and their encoders (0.5 hr). It was photometric with decent seeing (0.6-0.8 arcsec) for the remainder of the night.

LBC:
  • Had a problem with PSF SX M1 collimation caused by PMCL running out of virtual memory. Restarting PMCL fixed it.
  • Had to roll back to the old version of track_first, new version failed.
  • DOFPIA did not converge for Blue -- it is doing something wrong with Z4/Z11/Z22 and oscillating badly until Z22 saturates and fills in the central hole.

MODS1:
  • Active optics and guiding was nominal - did not take any MODS1 science images
  • Have not done any MODS1 AGw alignment yet (adjustment to the rotator center).
  • We took a set of field aberration data on BS9107 at the end of the night.

LUCI1:
  • Active optics and guiding was nominal - did not take any LUCI1 science images
  • Have not done any LUCI1 AGw alignment yet. M3 seems obviously (but mildly) out of position by comparing LUCI1 pointing to MODS1.

Wake-up for telescope, LBCs, MODS1 and LUCI1 has been completed. Starting the beginning of SX M2 alignment. And we did in 1 night what it took 2 weeks to accomplish last year!!!

Nominal plan for tomorrow night: Align SX M3 and begin the AGW1 precision alignment. LUCI1@LFBG

Preparations

Details

02:00 Sky is mostly clear, but we are waiting for the humidity to go down.

Test Presets and Authorizations

02:08 Geno's LBC BOTH preset

02:09 Turn on LBC

02:14 BOTH preset from LBC WT10_300

02:24 Authorize MODS1, but this hangs at the homing the stages step Had to kill GCSL to escape..........

02:28 Open both vent doors midway

02:38 Trying to turn on MODS1 AGw

mods1data:lbto% isisCmd m1.ie calib reset
*** ERROR: CALIB CALIB=TIMEOUT cannot write to 192.168.139.111:8001

I think this means the Comtrol units need reset. Dave T. was working on it.

02:41 LUCI authorization succeeded

02:45 Move to horizon to wake-up DX AdSec (Runa)

03:17 Back to LBC authorization on both sides

03:20 All vent doors full open. Restart Mirror Ventilation

Open

03:27 Open shutter doors ---> They unlatched but did not open. Called DCox (They couldn't test the encoders in the rain.)

03:52 Doors are opening......... (%Open indicator is not working, but limit switches are OK)

03:57 Open and latched!

03:58 LBC preset tp WT10_300 (AZ=259, EL=70)

04:04 DOFPIA struggling to converge in poor seeing. Telescope is cooling off, but still ~4 degC above ambient. Blue seems to be running away in negative spherical.

04:15 Clear Active Optics on SX

04:16 DOFPIA again ---> Now the SX Z11 is better, but it is actually Z22 running away. Perhaps there are issues from the astigmatism and trefoil caused by actuator 410.

Issue 4778 for SX Primary Actuator 410.

04:26 Adjust SX Z22 manually......

04:29 Z22 Instrument Offset Gain to 0.0 --- but I got fooled by the GUI again

04:30 Clear Active Optics on SX

04:31 DOFPIA (stopped after 1 iteration)

04:34 DOFPIA

04:41 Instrument Offset Gain for Z22 really to zero now.

04:46 DOFPIA Blue has a jump in focus, and back of 7400 nm Z4. Why?

04:48 SX "Main Mirror Failed to focus"

Pressing "Collimate" on PSF says "command thread creation failed"

04:50 Stopping and restartig PSFL doesn't help.

04:56 Stop PSFL and restart on TCS4 (has been on TCS3)

Couldn't reach Dan C, called Chris B.

05:06 Restart both PMCs - PMCL had run out of virtual memory. That fixed it.

05:10 LBC Preset to WT10_274

05:15 DOFPIA

05:21 DOFPIA, /FIRST

05:27 Stop DOFPIA manually

05:28 WT10_274 x120 - guiding fails on both cameras

RPC Client/Host communication failed retval:5 [src/trackers/trackers.c:1653]

05:31 - second try also fails

05:37 Turn off LBC (it crashed on power down); lbckill/lbcstart; Power on

05:45 resend preset WT10_274

05:47 WT10_274 x120 - guiding fails on both cameras, but this time red is different, hanging at 32 sec working (blue was out of focus)

05:48:42 "image data is zero" message in LBC log, but images have stars in them.

05:56 DOFPIA - blue still oscillating in focus

06:07 WT10_274 x120

06:12 Call Kellee

06:18 Replace track_first.exe with the vintage version track_first.20081008.exe (same for track_next)

06:20 Power off LBC - it reports a crash. Turn On - it reports another crash.

06:25 WT10_274 x120 - now guiding is working - so it seems that Kellee's track_first executable was bad. Good images on red, but still ugly 2.5 arcsec images on blue.

06:30 Clear Active Optics on SX Primary and put gains back to nominal.

06:31 DOFPIA, /FIRST - Both red and blue pupils look better now. Blue still is going through large oscillations in Z4/Z11/Z22, so it isn't converging on the correct focus, even though it was not far off from the raw collimation model. Eventually is drives to Z22=-300, and by then the central hole is filled in.

06:44 WT10_274 x240 - this time there are terrible pupils in the guide images. Did the backout fail when I stopped DOFPIA? The last extrafocal images were OK. LBC seems to be exposing for 3840 sec even though I asked for a scaling of 240 --- Oops this is the rVfastextra ob..

The initial thumbnail is still stale, but then it changes to a question mark (correctly?)

06:50 WT10_274 x240

06:55 Add +200 to SX Z22 (net -198), but maybe that isn't the issue. Note that Red has +913 Z11 paired with -379 Z22, while Blue has -1500 Z11 paired with -397 Z22.

06:56 WT10_274 x240

LBC is now working OK, except DOFPIA isn't able to focus Blue. Has something about Blue changed to cause this? Or is it the same old problem?

MODS1

07:01 Switch over to MODS1

07:18 MODS1 ACQUIRE preset - pointing was only off by ~20 arcsec.

07:23 active acquisition fails on defocussed star. Then Geno wiped out the global offsets, so it disappeared altogether.

07:30 Active Optics is converging OK despite lots of initial coma and astigmatism. We need to adjust the default collimation later when the telescope structure is in better equilibrium. Guider image is 0.8 arcsec FWHM at EL=58.

Why is the guider magnitude dropping when active optics updates? See the guide plot between 7.50 and 7.65 UT. The answer is that GCS is shortening the exposure time.

07:47 offset 20 0 (RADEC ABS)

Guide RMS is 0.08 arcsec RMS, but the Y-direction has an occasional blip of ~0.3 arcsec. Are those atmospheric? or telescope-induced?

Geno can't get a faceplace temperature plot from PMCL MATLAB, but I can get one remotely from obs1.

LUCI1

08:00 Switch over to LUCI1

08:14 execMODS modsSleep.pro to put MODS to sleep

08:17 LUCI1 ACQUIRE preset - did not find a star

08:22 found it farther away on an M5 star -- M3 must be slightly out of position.

08:28 LUCI1 ACTIVE preset - converges to guider image of 0.55 arcsec FWHM (0.62 on wfs)

08:39 offset 20 0

MODS1

08:48 Switch back to MODS1 for Remounted M2 alignment, but wait .... the SX M2 was not dismounted during shutdown. It was dismounted in late spring and Doug realigned it on 20130521 UT. Am I remembering that correctly?

See BGTechProcedures for the procedure to realign M2 hexapod after the secondary unit has been remounted.

09:04 John forgot to open the MODS1 hatch, but after doing that we still don't see a star. I had to do a calib reset before the hatch would open?

09:15 Active Preset to BS9101

mkdir lbt/work/lbto/20130901_mods1

cd ........

agw_set_station, "LDG"
iif_register
field_collect

09:23 Preset to BS9107 at EL=56.6 as there wasn't a field list for BS9101. Seeing on guider is 0.9-1.1, wfs says 0.76.

We have NOT recentered the MODS1 agw on the rotator center.

09:27 field_collect with BS9107_field.list

failed at fifth star gs=50 (too far out?)

09:53 field_collect with BS9107_new_field.list

failed at star 50 again......

Doug M: Please have a look at this data when you get to the mountain. Based on the numbers scrolling by, binodal is not horrible.

10:37 field_collect, start=10 with BS9107_new_field.list

10:41 field_collect, start=11 with BS9107_new_field.list

Is there charge transfer smearing in the X-direction on the MODS1 guider? I'm getting the impression of a bias in ellipticity.

Also note that the off-axis presets are systematically missing the final offset in the Y direction. The initial guide corrections are not random.

GStar wfsc #
0 95-98
43 99-101
0 102-103
35 104-106
0 107-108
44 109-111
0 112
36 113-115
0 116-117
GStar wfsc #
0 118-119
23 120-122
0 123
28 124-126
0 127
36 128-130
0 131-132
33 133-135
0 137-138
18 139-141
0 142
19 143-145
0 146
24 147-149
0 150
42 151-153
0 154
20 155-157
0 158-159
0 160
39 163-165
0 166
4 167-169
0 170
43 171-173
0 174
44 175-177
0 178
12 179-181
0 182
41 183-185
0 186
22 187-189
0 190
6 191-193
0 194
9 195-197
0 198

20130903 Fiel data analysis DLM
  • Result show no binodal within noise of measurements
  • No M2 position change

 => New Field Data Results
             Measured bi-nodal astigmatism
              From Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  -0.5", Delta IE =   0.5"
         Tilt =   1.9", Delta CA =  -1.9"
  
                M2 Mis-alignment 
             M2 Arcseconds and Millimeters
         -------------------------------------------
          Delta RX =    -5.", Delta Y = -0.025 mm
          Delta RY =   -19.", Delta X = -0.098 mm

Close

11:23 Close for impending dawn - the shutter doors were happy to close - thanks Dan.

-- JohnHill - 01 Sep 2013
  • Mods1 Field Astigmatism:
    mods_field.jpg
Topic attachments
I Attachment ActionSorted ascending Size Date Who Comment
mods_field.jpgjpg mods_field.jpg manage 27 K 03 Sep 2013 - 07:04 DougMiller Mods1 Field Astigmatism
Topic revision: r3 - 03 Sep 2013, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback